### v2

@@ 5.2. Benchmark Metrics on Runtime
The execution time of the $\Delta$-Motif algorithm comprises two phases: data preparation and computation. In the data preparation phase, the embeddings of required motif set, $Res(M)$, are computed and stored as dataframes for reuse. The computation phase then decomposes the pattern graph using VF2 and executes dataframe join-and-filter operations to produce the final matches.
In many scenarios, the same data graph is queried repeatedly with different pattern graphs, such as in quantum circuit compilation (Nation2023, 9) or in streaming subgraph isomorphism (Duong2021, 50), which arises in domains including social network analysis, chemistry, and biology. In such cases, the data preparation phase can be performed once, cached, and reused, allowing its cost to be excluded from subsequent queries and leaving only the pattern-searching phase. This reuse can deliver significant runtime savings and further speedups. To highlight this potential, we also report results with data preparation time omitted, in Section 6 marked as $\Delta$-Motif GPU.
@@ 6. Benchmarks and Results
Table 1. Data science relevant graphs used for benchmarking, all taken from the SuiteSparse Matrix Collection (ssmc, 51). ID denotes the SSMC identifier. All matrices are square and the dimension refers to both rows and columns. Non-zero entries refers to the total number of non-zero entries.
We evaluate our GPU-accelerated, motif-based subgraph isomorphism approach using two complementary benchmark suites. The first measures the performance of $\Delta$-Motif on large data graphs with small pattern graphs, such as social network analysis workloads, as well as on fixed data graphs with scaling pattern sizes. The second examines its practical impact when integrated into an end-to-end quantum layout selection pipeline.
6.1. Social Network Analysis Using Real World Networks
Triangle-counting, clustering coefficients, and $k$-truss analysis are all based on finding and enumerating triangles ($M_{3\text{-}O}$) in a graph and are extremely popular metrics in the fields of social network analysis, cyber-security and biological networks. Given its significance, our first benchmark evaluates $\Delta$-Motif using $M_{3\text{-}O}$ as the pattern graph on several datasets from the SparseSuite Matrix Collection (ssmc, 51). Table 1 summarizes these tested networks. For instance, coAuthorsCiteseer and citationCiteseer are representative social networks with irregular topologies that follow a power-law distribution. We compare our $\Delta$-Motif algorithm on GPU against VF2 algorithm on CPU, and report both the wall-clock times and the speedups in Fig. 3.
[…]
Table 2. Range of observed speedups for $\Delta$-Motif (GPU) over VF2 for each pattern size and data graph configuration. For every configuration, the table lists the mean, minimum, and maximum speedups measured across all random seeds.
@@ 6.1. Social Network Analysis Using Real World Networks
[…]
@@ 6.2. Subgraph Enumeration With Synthetic Subgraphs
Table 2 confirms a consistent performance advantage for $\Delta$-Motif GPU across all configurations, with mean speedups ranging from $3\times$ on small patterns (which typically under-utilize the GPU) and up to $53.43\times$ on the largest cases. Larger data graphs generally yield higher speedups; for example, increasing the square grid size from 1600 to 3600 vertices raises the mean speedup for 100-node patterns from $15.20\times$ to $53.43\times$. Medium-to-large patterns (60 nodes and above) generally achieve the highest speedups, particularly for the 3600-node square grid and 4485-node heavy-hex, where the parallel join-and-filter phase is fully leveraged, reaching maximum observed speedups of $415.20\times$ and $594.92\times$, respectively. While variability is observed across seeds, the minimum speedup remains above the VF2 baseline in nearly all cases. The few low-minimum cases (e.g., heavy-hex with 20-node patterns) occur when the pattern size is small enough that GPU launch overheads dominate the runtime. Overall, the results show that $\Delta$-Motif consistently outperforms VF2, with the largest benefits realized on large data graphs and medium-to-large pattern sizes that maximize GPU utilization. We also note that pattern graph decomposition accounts for less than 1% of the total runtime in all cases, with overall execution completing in just a few seconds.
Figure 4. Performance comparison on two data graph topologies: (a)–(c) heavy-hex lattice (4,485 nodes) and (d)–(f) 2D square grid (3,600 nodes), using 50 random 60-node pattern graphs. Panels show (a)(d) total execution time, (b)(e) speedup relative to VF2, and (c)(f) number of decomposition iterations. Bars represent mean values, with error bars indicating the minimum and maximum. Results are reported for VF2 and $\Delta$-Motif on CPU and GPU, as well as $\Delta$-Motif GPU with data preparation excluded.
@@ 6.5. End-to-End Transpilation Benchmark
In this benchmark, we integrated the $\Delta$-Motif algorithm into a complete quantum transpilation pipeline for layout selection. We compared our approach against Qiskit’s VF2PostLayout transpiler pass (qiskit2024, 49), which relies on rustworkx’s VF2 implementation for layout generation, followed by a sequential Python-based scoring function for candidate scoring (Nation2023, 9). [^2] The scoring stage is sequential for undirected data graphs, which we use because they yield more isomorphisms. While some quantum hardware supports only directed two-qubit operations, reversing the direction in the postprocessing is trivial in practice, allowing us to consider a much larger space of feasible layouts using the undirected model. Note that although the design of scoring cost functions is an active research topic in quantum computing, it is not the focus of this work. Here, we adopt the same cost function as used in (Nation2023, 9).
The data graph was derived from the qubit connectivity of IBM’s ibmq_fez, a state-of-the-art 156-qubit quantum processor (represented as a 156-node graph). We benchmark our pipeline on a diverse set of pattern graphs obtained from quantum circuits in the QASMBench benchmark suite (qasmbench, 52), covering sizes from 30 to 70 qubits and including a scaling set of Bernstein–Vazirani (BV) circuits. [^3] Before layout selection, all input logical circuits are pre-transpiled with Qiskit (opt_level=1) to ensure that every two-qubit operation is mapped to physically connected qubits, making the resulting pattern graphs valid subgraphs of the device connectivity.
@@ 7. Conclusion
We implement $\Delta$-Motif using open-source, commodity software that effectively harnesses the parallelism of NVIDIA GPUs. Our design avoids hardware-specific kernels, instead capitalizing on the highly optimized RAPIDS libraries, making the method portable across accelerator hardware. This implementation routinely achieves order-of-magnitude improvements over VF2, with observed speedups of up to $595\times$ in optimal cases. To the best of our knowledge, $\Delta$-Motif is the first GPU-based solution for subgraph isomorphism built entirely on database primitives. Given the uniqueness of our solution, we believe this approach not only provides a powerful new tool for graph analysis but also opens new directions for research at the intersection of databases, graph algorithms, and high-performance computing.

### v3

@@ 5.2. Benchmark Metrics on Runtime
The execution time of the $\Delta$-Motif algorithm comprises two phases: data preparation and computation. In the data preparation phase, the embeddings of required motif set, $Res(M)$, are computed and stored as dataframes for reuse. The computation phase then decomposes the pattern graph and executes dataframe join-and-filter operations to produce the final matches.
In many scenarios, the same data graph is queried repeatedly with different pattern graphs, such as in quantum circuit compilation (undefh, ) or in streaming subgraph isomorphism (undefaat, ), which arises in domains including social network analysis, chemistry, and biology. In such cases, the data preparation phase can be performed once, cached, and reused, allowing its cost to be excluded from subsequent queries and leaving only the pattern-searching phase. This reuse can deliver significant runtime savings and further speedups. To highlight this potential, we also report results with data preparation time omitted, in Section 6 marked as $\Delta$-Motif GPU.
@@ 6. Benchmarks and Results
Table 1. Data science relevant graphs used for benchmarking, all taken from the SuiteSparse Matrix Collection (undefaau, ). ID denotes the SSMC identifier. All matrices are square and the dimension refers to both rows and columns. Non-zero entries refers to the total number of non-zero entries.
Table 2. Observed speedup ranges of $\Delta$-Motif (GPU) over VF2 across pattern sizes and data graph configurations. For each configuration, we report the mean, minimum, and maximum speedups measured over all random seeds. GSI performance is discussed in the main text but omitted from this table, as results were unavailable for most problem sizes.
@@
We evaluate our GPU-accelerated, motif-based subgraph isomorphism approach using two complementary benchmark suites. The first measures the performance of $\Delta$-Motif on large data graphs with small pattern graphs, such as social network analysis workloads. The second evaluates performance on graph instances motivated by quantum circuit compilation, which feature structured data graphs and comparatively large pattern graphs.
6.1. Social Network Analysis Using Real World Networks
Figure 3. Wall-clock runtimes and speedups for enumerating all $M_{3\text{-}O}$ subgraphs. The x-axis lists the target graphs (see Table 1). Panel (a) reports end-to-end runtimes (data preparation + computation) for $\Delta$-Motif on GPU, GSI and VF2 on CPU. Panel (b) shows the corresponding speedups, computed as $T_{\text{VF2}}/T_{\text{GSI}}$ and $T_{\text{VF2}}/T_{\Delta\text{-Motif}}$.
Triangle-counting, clustering coefficients, and $k$-truss analysis are all based on finding and enumerating triangles ($M_{3\text{-}O}$) in a graph and are extremely popular metrics in the fields of social network analysis, cyber-security and biological networks. Given its significance, our first benchmark evaluates $\Delta$-Motif using $M_{3\text{-}O}$ as the pattern graph on several datasets from the SparseSuite Matrix Collection (undefaau, ). Table 1 summarizes these tested networks. For instance, coAuthorsCiteseer and citationCiteseer are representative social networks with irregular topologies that follow a power-law distribution. We compare our $\Delta$-Motif algorithm on GPU against VF2 algorithm on CPU and GSI on GPU, and report both the wall-clock times and the speedups in Fig. 3. We also evaluated EGSM on these datasets, but it produced inconsistent results on loop-like pattern graphs, so we do not report its performance in the benchmarks.
For most of the tested networks, VF2 requires well over 200 seconds and can take up to 385 seconds in some cases. In contrast, GSI completes within 2–10 seconds, while $\Delta$-Motif on GPU consistently finishes in just 1–2 seconds. This substantial reduction in runtime demonstrates the ability of $\Delta$-Motif to efficiently handle real-world social network graphs of varying size and complexity. Overall, $\Delta$-Motif achieves speedups of $82\times$$323\times$ on a single GPU compared to a single CPU core running VF2, whereas GSI achieves speedups of $28\times$$145\times$.
Such improvements are particularly important for application domains such as social network analysis and cyber-security, where subgraph matching is often performed repeatedly at scale under strict time constraints. As a point of reference, achieving comparable performance using CPUs alone would require parallel execution across approximately 82–323 cores with near-perfect scalability, which is difficult to attain in practice given the limited number of cores within a single socket. By contrast, $\Delta$-Motif leverages a data-centric, massively parallel formulation that maps naturally to GPU hardware, making large-scale subgraph analysis both faster and more practical.
@@ 6.2. Subgraph Enumeration With Synthetic Subgraphs
Table 2 confirms a consistent performance advantage for $\Delta$-Motif GPU across all configurations, with mean speedups ranging from $3\times$ on small patterns (which typically under-utilize the GPU) and up to $53.43\times$ on the largest cases. For smaller patterns, $\Delta$-Motif under-utilizes the GPU, and fixed launch and data-movement overheads constitute a larger fraction of the total runtime. Larger data graphs generally yield higher speedups; for example, increasing the square grid size from 1600 to 3600 vertices raises the mean speedup for 100-node patterns from $15.20\times$ to $53.43\times$. Medium-to-large patterns (60 nodes and above) generally achieve the highest speedups, particularly for the 3600-node square grid and 4485-node heavy-hex, where the parallel join-and-filter phase is fully leveraged, reaching maximum observed speedups of $415.20\times$ and $594.92\times$, respectively. While variability is observed across seeds, the minimum speedup remains above the VF2 baseline in nearly all cases. The few low-minimum cases (e.g., heavy-hex with 20-node patterns) arise when the pattern size is sufficiently small that GPU launch overhead dominates the runtime.
We additionally compare $\Delta$-Motif against GSI where results are available. On smaller instances, such as the 1600-node square grid with 20-node patterns, the two methods show comparable average performance, measured by the mean speedup $T_{\text{GSI}}/T_{\Delta\text{-}Motif}$, with maximum speedups of up to $17\times$. As problem difficulty increases, the performance gap widens. For the 3600-node square grid with 20-node patterns, $\Delta$-Motif is on average $2\times$ faster than GSI, with maximum speedups reaching $63\times$. For larger patterns, GSI fails to complete within a reasonable time window, while $\Delta$-Motif continues to scale effectively.
Overall, the results show that $\Delta$-Motif consistently outperforms VF2 and increasingly outperforms GSI as problem size and complexity grow, with the largest benefits realized on large data graphs and medium-to-large pattern sizes that maximize GPU utilization. We also note that pattern graph decomposition accounts for less than 1% of the total runtime in all cases, with end-to-end execution completing in just a few seconds.
Figure 4. Performance comparison on two data graph topologies: (a)–(c) heavy-hex lattice (4,485 nodes) and (d)–(f) 2D square grid (3,600 nodes), using 50 random 60-node pattern graphs. Panels show (a)(d) total execution time, (b)(e) speedup relative to VF2, and (c)(f) number of decomposition iterations. Bars represent mean values, with error bars indicating the minimum and maximum. Results are reported for VF2 and $\Delta$-Motif on CPU and GPU, as well as $\Delta$-Motif GPU with data preparation excluded. GSI did not complete when the pattern graph had 60 vertices.
@@ 6.5. End-to-End Transpilation Benchmark
In this benchmark, we integrated the $\Delta$-Motif algorithm into a complete quantum transpilation pipeline for layout selection. We compared our approach against Qiskit’s VF2PostLayout transpiler pass (undefaar, ), which relies on rustworkx’s VF2 implementation for layout generation, followed by a sequential Python-based scoring function for candidate scoring (undefh, ). [^2] The scoring stage is sequential for undirected data graphs, which we use because they yield more isomorphisms. While some quantum hardware supports only directed two-qubit operations, reversing the direction in the postprocessing is trivial in practice, allowing us to consider a much larger space of feasible layouts using the undirected model. Note that although the design of scoring cost functions is an active research topic in quantum computing, it is not the focus of this work. Here, we adopt the same cost function as used in (undefh, ).
The data graph was derived from the qubit connectivity of IBM’s ibmq_fez, a state-of-the-art 156-qubit quantum processor (represented as a 156-node graph). We benchmark our pipeline on a diverse set of pattern graphs obtained from quantum circuits in the QASMBench benchmark suite (undefaav, ), covering sizes from 30 to 70 qubits and including a scaling set of Bernstein–Vazirani (BV) circuits. [^3] Before layout selection, all input logical circuits are pre-transpiled with Qiskit (opt_level=1) to ensure that every two-qubit operation is mapped to physically connected qubits, making the resulting pattern graphs valid subgraphs of the device connectivity.
@@ 7. Conclusion
We implement $\Delta$-Motif using open-source, commodity software that effectively exploits the parallelism of NVIDIA GPUs. Our design avoids hardware-specific kernels and instead leverages highly optimized RAPIDS libraries, enabling portability across accelerator platforms. This implementation consistently achieves order-of-magnitude performance improvements over VF2, with observed speedups of up to $595\times$ in favorable cases, and outperforms the GPU-based approach GSI in benchmarks where direct comparison is applicable. To the best of our knowledge, $\Delta$-Motif is the first GPU-based subgraph isomorphism approach implemented entirely using database-style relational primitives. Given the uniqueness of our solution, we believe this approach not only provides a powerful new tool for graph analysis but also opens new directions for research at the intersection of databases, graph algorithms, and high-performance computing.